OpenStudy (anonymous):

1. Use Calc. to determine the critical points, local extrema, inflection points, and intervals where f(x) is concave up or down. f(x) = x^4 + 2x^3 - 1

OpenStudy (jack1):

whoa... so u just want...everything... then...?

OpenStudy (jack1):

OK @amelapal ... so first of all, what are the turning points...?

OpenStudy (anonymous):

well the extrema is the hardest part for me I don't know why :-(

OpenStudy (anonymous):

okay so the critical points are when the function is equal to zero right?

OpenStudy (anonymous):

**the function's derivative

OpenStudy (jack1):

sorry, hit send and saw ur post, my bad

OpenStudy (anonymous):

its alright just caught my mistake!

OpenStudy (anonymous):

so in this case the derivative would 4x^3 + 6x^2?

OpenStudy (jack1):

perfect

OpenStudy (anonymous):

so one critical point can be zero right?

OpenStudy (jack1):

yep, but sub back in to initial eqn prove

OpenStudy (jack1):

*to prove

OpenStudy (anonymous):

plug zero back into the original function? it gives you negative one, what do I do with that?

OpenStudy (anonymous):

@Jack1

zepdrix (zepdrix):

Sooo you've got zero, any others? :)

OpenStudy (anonymous):

No! For some reason I can't find any other numbers!

zepdrix (zepdrix):

\[\Large\bf\sf 0=4x^3+6x^2\]Factoring an x^2 out of each,\[\Large\bf\sf 0=x^2(4x+6)\] x^2=0 gives us our critical point at x=0. How bout the other factor?

OpenStudy (anonymous):

4x + 6 = 0 4x = -6 x= -6/4 ?

zepdrix (zepdrix):

Ok good, which simplifies to x=-3/2 or x=-1.5

zepdrix (zepdrix):

Those are the critical points of the function. So what's next? Determining max and min?

OpenStudy (anonymous):

that would be another critical point?

OpenStudy (anonymous):

wait I get it! okay yes finding max and min now

OpenStudy (anonymous):

you plug both 0 in the original function and get -1 and plug -6/4 and get -2.6875

zepdrix (zepdrix):

No we don't want to do that! :Oo

zepdrix (zepdrix):

So like ummm

OpenStudy (anonymous):

:o

zepdrix (zepdrix):

`critical points` / `stationary points` / `equilibrium points` they go by many different names, depending on what field of work you're in. It's a special place where the function is not `increasing` nor is it `decreasing`. The function is at the bottom of a valley, or at the top of a hill. So what we want to do is, check out what's happening on each side of the critical point.

zepdrix (zepdrix):



zepdrix (zepdrix):

Here is a nice little shortcut, We only care about the `sign` of the result. \[\Large\bf\sf f'(x)=x^2(4x+6)\]See how we have a square that we factored out? That will ALWAYS be positive, no matter what we plug into the function. So we can ignore that part.\[\Large\bf\sf f'(x)=(positive)(4x+6)\]Let's plug -2 into the derivative function.

zepdrix (zepdrix):

\[\Large\bf\sf f'(-2)=(positive)(-8+6)\]\[\Large\bf\sf f'(-2)=(positive)(negative)\]\[\Large\bf\sf f'(-2)=negative\]

zepdrix (zepdrix):

Negative tells us that the function is `decreasing` on the left side of x=-1.5

zepdrix (zepdrix):



zepdrix (zepdrix):

Does this process make sense .. a lil bit maybe? :o

zepdrix (zepdrix):

We have 2 more areas we have to check. We want to know what's happening on the right side of x=-1.5 and then on the right side of x=0.

zepdrix (zepdrix):

brb i need some chocolate milk +_+

zepdrix (zepdrix):

You look at the stuff, think :U lemme know if it's too crazy.

OpenStudy (anonymous):

it makes total sense! but brb I have to get ready for school now -_-!

zepdrix (zepdrix):

oh heh :3

OpenStudy (jack1):

apologies @amelapal , laptop issues, @zepdrix finished it up though